# Equilibrium under Time-Inconsistency: A New Existence Theory by Vanishing Entropy Regularization

### 11 Mar 2026 v1

@@ 3 Time-inconsistent Control under Entropy Regularization
Suppose that Assumption 3.1 holds. There exists a $\lambda_{0}>0$ such that the following estimate holds for all $0<\lambda\leq\lambda_{0}$,
$|\mathcal{H}(\Gamma_{\lambda}(x,y,a))|=\left|\int_{U}\ln(\Gamma_{\lambda}(x,y,a))\Gamma_{\lambda}(x,y,a)da\right|\leq K_{6}+K_{7}|\ln\lambda|+K_{8}\ln(1+|y|),\quad\forall y\in\mathbb{R}^{d},$
where $K_{6},K_{7},K_{8}$ are positive constants that depend on $\ell$, $\operatorname{\texttt{Leb}}(U)$ and the constants $\Theta$, $\zeta$ and $\gamma$ in Assumption 3.1, but they are independent of $\lambda$.
For a given $w(x)\in{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})$, we next provide Hölder norm estimates for certain terms that appear in the policy
@@
Suppose Assumptions 2.1, 2.2 hold. Given $w(x)\in{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})$ such that $\|w\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq M$ for a finite constant $M\geq 1$, the following estimates hold for the policy $\pi$ in (3.12) that
$\|\lambda\mathcal{H}(\pi)\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})}\leq A_{0}M\exp(\frac{A_{0}M}{\lambda}),$
$\|b^{\pi}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})},\|r^{\pi}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq\frac{A_{0}M}{\lambda}\exp(\frac{A_{0}M}{\lambda}),$
where $A_{0}$ is a constant depending on $d$, $\operatorname{\texttt{Leb}}(U)$ and the parameters $K_{1},K_{3}$ in Assumptions 2.1 and 2.2, but independent of $\lambda\leq\lambda_{0}$ and $M$.
@@
Throughout this proof, let $A_{0}$ be a generic positive constant depending on $d$, $\operatorname{\texttt{Leb}}(U)$ and $K_{0},K_{1},K_{3}$ in Assumptions 2.1 and 2.2 but independent of $\lambda$ and $M$, which may vary from line to line.
Let $g(t,x,a):=\frac{1}{\lambda}[b(x,a)D_{x}w(x)+r(0,x,a)]$, a direct calculation gives that
@@
In view of (3.15), the first term on the right-hand side above is estimated by
@@
$=\left|g(x,a)-g(y,a)-\ln\left(\int_{U}\exp[g(x,a^{\prime})-g(y,a^{\prime})]da^{\prime}\right)\right|\leq\frac{A_{0}M}{\lambda}|x-y|^{\alpha}.$
@@
$\left|\frac{\pi(y,a)}{\pi(x,a)}-1\right|\leq$ $\frac{\exp[g(y,a)]\int_{U}\exp\left[g(y,a^{\prime})+[g(\cdot,a)]_{\alpha/2,\alpha}|x-y|^{\alpha}\right]da^{\prime}}{\int_{U}\exp[g(y,a^{\prime})]da^{\prime}\exp[g(x,a)]}-1$
@@
$\leq$ $\frac{A_{0}M}{\lambda}\exp(\frac{A_{0}M}{\lambda})|x-y|^{\alpha},$
for all $|x-y|^{\alpha}\leq 1$. Combining above with (3.15) and (3.18), we conclude that
@@
Let Assumptions 2.1, 2.2, 3.1, and conditions (3.2)–(3.3) hold. For a given $w(x)\in{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})$, consider the feedback control
@@
Then $V^{\pi}_{\lambda}(t,x)$ defined in (3.1) satisfies the following PDE
$\partial_{t}V^{\pi}(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}V^{\pi}(t,x)\right)+b^{\pi}(x)D_{x}V^{\pi}(t,x)+r^{\pi}(t,x)+\lambda\delta(t)\mathcal{H}(\pi(x))=0.$ (3.19)
Moreover, there exists a constant $A^{*}>0$ (resp. $A_{\lambda}$) depending on $d,\ell$, $\alpha$, $\operatorname{\texttt{Leb}}(U)$ and the parameters $K_{0},\cdots,K_{5}$, $\eta$, $\Theta,\zeta,\gamma$ in Assumptions 2.1, 2.2, 3.1, and conditions (3.2)–(3.3), but independent of $\lambda\leq\lambda_{0}$, (resp. further depends on $\lambda$) such that if $\|w\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq A^{*}$, then
$\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\vee\|V^{\pi}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}\vee\sup_{x\in\mathbb{R}^{d}}\|V^{\pi}_{\lambda}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*};$
$(\text{resp. }\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A_{\lambda}).$
@@
Fix $0<\lambda\leq\lambda_{0}$. Take $w\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, and set $\pi(x,a):=\Gamma_{\lambda}(x,D_{x}w(0,x))$. Combing Assumptions 2.1, 2.2, 3.1 and (3.2), we get that
$\|b^{\pi}\|_{{\mathcal{C}}^{0}_{\alpha}(\mathbb{R}^{d})},\;\|r^{\pi}\|_{{\mathcal{C}}^{0,0}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})},\;\|\delta\lambda\mathcal{H}(\pi)\|_{{\mathcal{C}}^{0,0}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}<\infty.$
Rewrite $b^{\pi(x)}(x)$, $r^{\pi(x)}(t,x)$ by $b^{\pi}(x)$, $r^{\pi}(t,x)$ respectively for short. It holds that the SDE
$dX^{\pi}_{t}=b^{\pi}(X^{\pi}_{t})dt+\sigma(X^{\pi}_{t})dW_{t}$
admits a unique strong solution. Itô’s formula yields that $V^{\pi}(t,x)$ satisfies the parabolic PDE (3.19). Moreover, by standard Schauder estimate, $\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}<\infty$.
We next analyze $\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}$ and $\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{1,1}(\mathbb{T}\times\mathbb{R}^{d})}$. Throughout the proof, let $A_{0}$ be a generic positive constant depending on $d,\alpha$ and $K_{1},\cdots,K_{5}$, $\eta$, $\Theta,\zeta,\gamma$ in Assumptions 2.1, 2.2, 3.1 and conditions (3.2)–(3.3), but independent of $\lambda$ and norms of all orders of derivatives of $w$, which may very from line to line.
By Lemma 3.1,
@@
Then for each $(t,x)\in\mathbb{T}\times\mathbb{R}^{d}$, it holds that
$\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0}(\mathbb{T}\times\mathbb{R}^{d})}\leq$ $\int_{-\infty}^{\infty}\left[\sup_{(x,a)\in\mathbb{R}^{d}\times U}|r(s,x,a)|+\delta(s)A_{0}\bigg(1+\ln\big(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\big)\bigg)\right]ds$
@@
$|\partial_{t}V^{\pi}_{\lambda}(t,x)|=$ $\mathbb{E}_{x}\left[\int_{-\infty}^{\infty}\left(\partial_{t}r^{\pi}(t+s,X^{\pi}_{s})+\delta_{t}(t+s)\lambda\mathcal{H}(\pi(X^{\pi}_{s}))\right)ds\right]$
$\leq$ $\int_{-\infty}^{\infty}\left[\sup_{(x,a)\in\mathbb{R}^{d}\times U}|\partial_{t}r(s,x,a)|+\delta_{t}(s)A_{0}\bigg(1+\ln(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\big)\bigg)\right]ds$
@@
where the inequalities above follow from (2.1) and (3.2). Because $\pi$ is independent of $t$, (2.2) and (3.3) together imply that, for $l=0,1$,
$\sup_{|t_{0}-t_{1}|\leq 1/2}\frac{\left|\partial^{l}_{t}V^{\pi}_{\lambda}(t_{0},x)-\partial^{l}_{t}V^{\pi}_{\lambda}(t_{1},x)\right|}{|t_{0}-t_{1}|^{\alpha/2}}$
$\leq\sup_{t_{1}\in[t_{0}-1/2,t_{0}+1/2]\cap\mathbb{T}}\mathbb{E}_{x}\int_{0}^{\infty}\frac{\bigg|\partial^{l}_{t}r(t_{0}+s,X^{\pi}_{s})-\partial^{l}_{t}r(t_{1}+s,X^{\pi}_{s})\bigg|}{|t_{0}-t_{1}|^{\alpha/2}}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\frac{\left|\partial^{l}_{t}\delta(t_{0}+s)-\partial^{l}_{t}\delta(t_{1}+s)\right|}{|t_{0}-t_{1}|^{\alpha/2}}\lambda\mathcal{H}(\pi(X^{\pi}_{s}))\bigg|ds$ (3.23)
$\leq\int_{0}^{\infty}\Bigg[\bigg(\sup_{(y,a)\in\mathbb{R}^{d}\times U,t_{1}\in[t_{0}-1/2,t_{0}+1/2]\cap\mathbb{T}}\frac{\left|\partial^{l}_{t}r(t_{0}+s,y,a)-\partial^{l}_{t}r(t_{1}+s,y,a)\right|}{|t_{0}-t_{1}|^{\alpha/2}}$
$\qquad\qquad\quad+\sup_{t_{1}\in[t_{0}-1/2,t_{0}+1/2]\cap\mathbb{T}}\frac{\left|\partial_{t}^{l}\delta(t_{0}+s)-\partial_{t}^{l}\delta(t_{1}+s)\right|}{|t_{0}-t_{1}|^{\alpha/2}}\bigg)A_{0}\bigg(1+\ln\big(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\big)\bigg)\Bigg]ds$
$\leq A_{0}\big(1+\ln\left(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\right)\big).$
Then, combining (3.21)–(3.23) leads to

### 04 Apr 2026 v2

@@ 3 Existence of Equilibrium under Entropy Regularization
Suppose that Assumption 3.1 holds. There exists a constant $\lambda_{0}>0$ such that the following estimate holds for all $0<\lambda\leq\lambda_{0}$,
$|\mathcal{H}(\Gamma_{\lambda}(x,p,a))|=\left|\int_{U}\ln(\Gamma_{\lambda}(x,p,a))\Gamma_{\lambda}(x,p,a)da\right|\leq K_{6}+K_{7}|\ln\lambda|+K_{8}\ln(1+|p|),\quad\forall x,p\in\mathbb{R}^{d},$
where $K_{6},K_{7},K_{8}$ are positive constants depending on $\ell$, $\operatorname{\texttt{Leb}}(U)$ and the constants $\Theta$, $\zeta$ and $\gamma$ in Assumption 3.1, but independent of $\lambda$.
The following lemma establishes Hölder norm estimates for key quantities induced by the policy
@@
Suppose Assumptions 2.1, 2.2 hold. If $w(x)\in{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})$ satisfies $\|w\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq M$ for a finite constant $M\geq 1$, then the following estimates hold for the policy $\pi$ in (3.12) :
$\|\lambda\mathcal{H}(\pi)\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})}\leq A_{0}M\exp(\frac{A_{0}M}{\lambda}),\quad\|b^{\pi}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})}\vee\|r^{\pi}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq\frac{A_{0}M}{\lambda}\exp(\frac{A_{0}M}{\lambda}),$
where $A_{0}$ is a constant depending on $d$, $\operatorname{\texttt{Leb}}(U)$ and the parameters $K_{0},K_{1},K_{3}$ in Assumptions 2.1 and 2.2, but independent of $\lambda\leq\lambda_{0}$ and $M$.
@@
Throughout this proof, let $|x-y|\leq 1$ and $A_{0}$ be a generic positive constant depending on $d$, $\operatorname{\texttt{Leb}}(U)$ and $K_{0},K_{1},K_{3}$ in Assumptions 2.1 and 2.2 but independent of $\lambda$ and $M$, which may vary from line to line.
Let $g(x,a):=\frac{1}{\lambda}[b(x,a)D_{x}w(x)+r(0,x,a)]$. A direct calculation gives that
@@
In view of (3.14)–(3.16), the first term on the right-hand side above is estimated by
@@
$\leq\left|g(x,a)-g(y,a)\right|+\left|\ln\left(\int_{U}\exp[g(x,a)]da\right)-\ln\left(\int_{U}\exp[g(y,a)]da\right)\right|\leq\frac{A_{0}M}{\lambda}|x-y|^{\alpha}.$
@@
$\left|\frac{\pi(y,a)}{\pi(x,a)}-1\right|\leq$ $\frac{\exp[g(y,a)]\int_{U}\exp\left[g(y,a^{\prime})+[g(\cdot,a)]_{\alpha}|x-y|^{\alpha}\right]da^{\prime}}{\int_{U}\exp[g(y,a^{\prime})]da^{\prime}\exp[g(x,a)]}-1$
@@
$\leq$ $\frac{A_{0}M}{\lambda}\exp(\frac{A_{0}M}{\lambda})|x-y|^{\alpha}.$
Combining above with (3.15) and (3.18), we conclude that
@@
Let Assumptions 2.1, 2.2, 3.1, and conditions (3.2)–(3.3) hold. Let $\lambda_{0}$ be given by Lemma 3.1 and fix $\lambda\in(0,\lambda_{0}]$. For a given $w(x)\in{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})$, consider the feedback control
@@
Then $V^{\pi}_{\lambda}(t,x)$ defined in (3.1) is the unique classic solution to the following PDE
$\partial_{t}u(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}u(t,x)\right)+b^{\pi}(x)D_{x}u(t,x)+r^{\pi}(t,x)+\lambda\delta(t)\mathcal{H}(\pi(x))=0,\,\,\forall(t,x)\in\mathbb{T}\times\mathbb{R}^{d}.$ (3.19)
And there exists a function $\Psi:\mathbb{T}\to\mathbb{T}$ and constants $A^{*},A_{\lambda}$, satisfying:
  • $\Psi(0)=1$, and $\Psi$ decreases with $\lim_{t\to\infty}\Psi(t)=0$,
  • $A^{*}$ depends only on $d,\ell$, $\alpha$, $\operatorname{\texttt{Leb}}(U)$ and the parameters $K_{0},\cdots,K_{5}$, $\eta$, $\Theta,\zeta,\gamma$ in Assumptions 2.1, 2.2, 3.1, and conditions (3.2)–(3.3), but independent of $\lambda\leq\lambda_{0}$; $A_{\lambda}$ further depends on $\lambda$,
such that if $\|w\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq A^{*}$, then the following estimate holds:
$\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}\vee\|V^{\pi}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}([t,\infty)\times\mathbb{R}^{d})}\leq A^{*}\Psi(t)\quad\forall t\in\mathbb{T};$ (3.20)
$\sup_{x\in\mathbb{R}^{d}}\|V^{\pi}_{\lambda}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*},\quad\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A_{\lambda}.$ (3.21)
@@
Throughout the proof, let $A_{0}$ be a generic positive constant depending on $d,\alpha$ and $K_{1},\cdots,K_{5}$, $\eta$, $\Theta,\zeta,\gamma$ in Assumptions 2.1, 2.2, 3.1 and conditions (3.2)–(3.3), but independent of $\lambda$ and norms of all orders of derivatives of $w$, which may vary from line to line.
Step 1. We first provide estimates on the solution to (3.19). Let $u(t,x)$ be a classic solution to (3.19). By Lemma 3.1,
@@
Define
$\Psi(t):=\frac{\sup_{x\in\mathbb{R}^{d},a\in U}\|r(\cdot,x,a)\|_{L^{\infty}([t,\infty))}\vee\|\delta\|_{L^{\infty}([t,\infty))}}{\sup_{x\in\mathbb{R}^{d},a\in U}\|r(\cdot,x,a)\|_{L^{\infty}(\mathbb{T})}\vee\|\delta\|_{L^{\infty}(\mathbb{T})}},$ (3.23)
then $\Psi(0)=1$, $\Psi(t)$ decreases, and we conclude from Assumption 2.1 and (3.2) that
$\lim_{t\to\infty}\Psi(t)=0.$ (3.24)
Recall that $p=\frac{d+2}{1-\alpha}$. For an arbitrary fixed $(t_{0},x_{0})\in\mathbb{T}\times\mathbb{R}^{d}$, Sobolev embedding (see, e.g., [19, II.3]) gives that
$\|u\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}\leq A_{0}\|u\|_{W^{1,2}_{p}(D_{2}(t_{0},x_{0}))}.$ (3.25)
$\|u\|_{W_{p}^{1,2}(D_{2}(t_{0},x_{0}))}$ $\leq A_{0}\bigg(\|r^{\pi}\|_{L^{\infty}(D_{3}(t_{0},x_{0}))}+\|\lambda\delta\mathcal{H}(\pi)\|_{L^{\infty}(D_{3}(t_{0},x_{0}))}\bigg).$
By combining this with (3.22) (3.23) and (3.25), we conclude for any $t\in\mathbb{T}$,
$\sup_{(t_{0},x_{0})\in[t,\infty)\times\mathbb{R}^{d}}\left(\|u\|_{W^{1,2}_{p}(D_{1}(t_{0},x_{0}))}\vee\|u\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}\right)$
$\leq A_{0}\left(\sup_{a\in U}\|r(\cdot,a)\|_{[t,\infty)\times\mathbb{R}^{d}}\vee\|\delta\|_{L^{\infty}([t,\infty))}\right)\left(1+\ln\left(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\right)\right).$
$\leq A_{0}\left(\sup_{a\in U}\|r(\cdot,a)\|_{\mathbb{T}\times\mathbb{R}^{d}}\vee\|\delta\|_{L^{\infty}(\mathbb{T})}\right)\Psi(t)\left(1+\ln\left(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\right)\right).$
This together with (2.1) and (3.2) yields, for any $t\in\mathbb{T}$ that,
$\|u\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}([t,\infty)\times\mathbb{R}^{d})}\vee\|u\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}\leq A_{0}\Psi(t)\left(1+\ln\left(1+\|w\|_{{\mathcal{C}}^{1}(\mathbb{R}^{d})}\right)\right).$ (3.26)
Step 2. Next, we show $V^{\pi}_{\lambda}$ is the unique classic solution to (3.19). Since $\|w\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}<\infty$, by Assumptions 2.1, 2.2, (3.2) and Lemma 3.2, we get that
$\|b^{\pi}\|_{{\mathcal{C}}^{0}_{\alpha}(\mathbb{R}^{d})},\;\|r^{\pi}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})},\;\|\delta\lambda\mathcal{H}(\pi)\|_{{\mathcal{C}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}<\infty.$
Then the SDE $dX^{\pi}_{t}=b^{\pi}(X^{\pi}_{t})dt+\sigma(X^{\pi}_{t})dW_{t}$ admits a unique strong solution. A standard application of Itô’s formula yields that $V^{\pi}_{\lambda}(t,x)$ satisfies the parabolic PDE (3.19). Now suppose $u$ is a classic solution to (3.19), and we show $V^{\pi}_{\lambda}=u$. Take an arbitrary $(t,x)\in\mathbb{T}\times\mathbb{R}^{d}$. Applying Itô- formula to $u(t+s,X^{\pi}_{s})$ yields
$du(t+s,X^{\pi}_{s})=$ $\left(\partial_{t}u(t+s,X^{\pi}_{s})+b^{\pi}(x)D_{x}u(t+s,X^{\pi}_{s})+\frac{1}{2}\operatorname{tr}\bigg((\sigma\sigma^{T})(X^{\pi}_{s})D^{2}_{x}u(t+s,X^{\pi}_{s})\bigg)\right)ds$
$+\sigma(X^{\pi}_{s})D_{x}u(X^{\pi}_{s})dW_{s}.$
Taking expectations on both sides leads to
$u(t,x)=$ $\mathbb{E}_{x}\bigg[\int_{0}^{T}\left(u_{t}(t+s,X^{\pi}_{s})+b^{\pi}(x)D_{x}u(t+s,X^{\pi}_{s})+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(X^{\pi}_{s})D^{2}_{x}u(t+s,X^{\pi}_{s})\right)\right)ds$
$\qquad+u(t+T,X^{\pi^{\infty}}_{T})\bigg]$
$=$ $\mathbb{E}_{x}\left[\int_{0}^{T}\left(r^{\pi}(t+s,X^{\pi}_{s})+\delta(t+s)\lambda\mathcal{H}(\pi(X^{\pi}_{s}))\right)ds\right]+\mathbb{E}_{x}\left[u(t+T,X^{\pi}_{T})\right]$
where the second equality follows from (3.19). By (3.24) and (3.26), $\lim_{T\to\infty}\ E_{x}\left[u(t+T,X^{\pi}_{T})\right]=0$. Therefore, by (2.1) in Assumption 2.1 and the Dominated Convergence Theorem, we conclude that the last line above converges to $\mathbb{E}_{x}\left[\int_{0}^{\infty}\left(r^{\pi}(t+s,X^{\pi}_{s})+\delta(t+s)\lambda\mathcal{H}(\pi(X^{\pi}_{s}))\right)ds\right]$. Consequently, it holds that $V^{\pi^{*}}_{\lambda}(t,x)=u(t,x)$ on $\mathbb{T}\times\mathbb{R}^{d}$.
Step 3. We now prove (3.20) and (3.21). By (3.1), for each $(t,x)\in\mathbb{T}\times\mathbb{R}^{d}$, it holds that
$\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0}(\mathbb{T}\times\mathbb{R}^{d})}\leq$ $\int_{0}^{\infty}\left[\sup_{(x,a)\in\mathbb{R}^{d}\times U}|r(s,x,a)|+\delta(s)A_{0}\bigg(1+\ln\big(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\big)\bigg)\right]ds$
@@
$|\partial_{t}V^{\pi}_{\lambda}(t,x)|=$ $\mathbb{E}_{x}\left[\int_{0}^{\infty}\left(\partial_{t}r^{\pi}(t+s,X^{\pi}_{s})+\delta_{t}(t+s)\lambda\mathcal{H}(\pi(X^{\pi}_{s}))\right)ds\right]$
$\leq$ $\int_{0}^{\infty}\left[\sup_{(x,a)\in\mathbb{R}^{d}\times U}|\partial_{t}r(s,x,a)|+\delta_{t}(s)A_{0}\bigg(1+\ln(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\big)\bigg)\right]ds$
@@
where the inequalities above follow from (2.1) and (3.2). Because $\pi$ is independent of $t$, we have, for $l=0,1$, that
$\sup_{t_{0}<t_{1}\leq t_{0}+1}\frac{\left|\partial^{l}_{t}V^{\pi}_{\lambda}(t_{0},x)-\partial^{l}_{t}V^{\pi}_{\lambda}(t_{1},x)\right|}{|t_{0}-t_{1}|^{\alpha/2}}$
$\leq\sup_{t_{0}<t_{1}\leq t_{0}+1}\mathbb{E}_{x}\int_{0}^{\infty}\Bigg(\frac{\big|\partial^{l}_{t}r(t_{0}+s,X^{\pi}_{s})-\partial^{l}_{t}r(t_{1}+s,X^{\pi}_{s})\big|}{|t_{0}-t_{1}|^{\alpha/2}}$
$\qquad\qquad\qquad\qquad\qquad+\frac{\left|\partial^{l}_{t}\delta(t_{0}+s)-\partial^{l}_{t}\delta(t_{1}+s)\right|}{|t_{0}-t_{1}|^{\alpha/2}}\lambda\mathcal{H}(\pi(X^{\pi}_{s}))\Bigg)ds$ (3.29)
$\leq\int_{0}^{\infty}\Bigg[\Bigg(\sup_{(y,a)\in\mathbb{R}^{d}\times U,t_{0}<t_{1}\leq t_{0}+1}\frac{\left|\partial^{l}_{t}r(t_{0}+s,y,a)-\partial^{l}_{t}r(t_{1}+s,y,a)\right|}{|t_{0}-t_{1}|^{\alpha/2}}$
$\qquad\qquad\quad+\sup_{t_{0}<t_{1}\leq t_{0}+1}\frac{\left|\partial_{t}^{l}\delta(t_{0}+s)-\partial_{t}^{l}\delta(t_{1}+s)\right|}{|t_{0}-t_{1}|^{\alpha/2}}\Bigg)A_{0}\bigg(1+\ln\big(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\big)\bigg)\Bigg]ds$
$\leq A_{0}\left(1+\ln\left(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\right)\right),$
where the last inequality follows from (2.2) and (3.3). Then, combining (3.27)–(3.28) leads to